Evaluate"\intop\intop" x2+y2 +2

Where Ris bounded by x2+ y2 =1


1
Expert's answer
2021-06-28T16:38:29-0400

Let us evaluate "\iint\limits_R(x^2+y^2+2)dxdy."


"\iint\limits_R(x^2+y^2+2)dxdy="


|"x=r\cos \varphi,\ y=r\sin\varphi,\ dxdy=rdrd\varphi, x^2+y^2=r^2" |


"=\int\limits_0^{2\pi}d\varphi\int\limits_0^1(r^2+2)rdr=2\pi\frac{1}{2}\int\limits_0^1(r^2+2)d(r^2+2)=
\pi\frac{(r^2+2)^2}{2}|_0^1=\pi(\frac{9}{2}-2)=\frac{5\pi}{2}."